OpenStudy (anonymous):

-x^2-2x+3=0 the solutions are????????

OpenStudy (anonymous):

-3,1

OpenStudy (anonymous):

use the quadratic formula http://en.wikipedia.org/wiki/Quadratic_formula#Quadratic_formula

OpenStudy (anonymous):

-x^2-2x+3=0 given x^2+2x-3=0 mutl by -1 (x+3)(x-1)=0 factor x+3=0 or x-1=0 theorem x=-3 or x=1

OpenStudy (anonymous):

or just factor like Mandolino did

jhonyy9 (jhonyy9):

divide by -1 x2+2x-3=0 (x+3)(x-1)=0 x=-3 x=1

OpenStudy (anonymous):

answer is x= -3 or, 1